Two trains on separate tracks move toward each other. Train 1 has a speed of 113 km/h; train 2, a speed of 59.0 km/h. Train 2 blows its horn, emitting a frequency of 500 Hz. What is the frequency heard by the engineer on train 1? HzA person wears a hearing aid that uniformly increases the intensity level of all audible frequencies of sound by 30.7 dB. The hearing aid picks up sound having a frequency of 250 Hz at an intensity of 2.70 x 101 W/m2. What is the intensity delivered to the eardrum? W/m2
Answer #1

Solution,

1)

Velocity, v1 = 113 km/h = 31.39 m/s

Velocity, v2 = 59 km/h = 16.39 m/s

Frequency, f' = f (u + v1) /(u - v2) = 500(340 + 31.39)/(340 - 16.39) = 573.82 Hz

Comment in case any doubt please rate my answer.....

According to the HOMEWORKLIB RULES I can answer one question at a time so please post other questions seperately.....